Shift report 20030309 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Mar 09 2003 - 15:56:11 EST

  • Next message: Apache: "Shift report 20030309 16:00-24:00"
    ------------------------------------------------------------
    Date: Sun Mar  9 15:56:10 2003, Shift: 08:00-16:00
    Supervisor: CEJ, with: RL
    Log book#14 pages 52 - 52
    
    Shift Summary:
    ------------------------------------------------------------
    1) Beam from 10:15 to 14:45. Quite low rates (ZDC starting 
       around 6000, ending around 1500)
    
    2) Ped, TPC ped and DVM runs: 8276, 8276 & 8274.
       MRS@90(A1050), FS@4(A1/1).
       8278 Trig 2 15K, trig 3 218K (D2 off the last 1/2 hour) 
       8281 Trig 2  4K, trig 3  53K (D2 on again)
    
    3) D2 went on stand-by fault mode at 11:57 (end of run 8278).
       No other problems reported.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Mar 09 2003 - 16:00:14 EST